son l is 17%, and on J it is 30%. Use decimals, not percents, in Hyac th Macaw invests 45% of her funds in stock l and the balance in stock J. The standard deviation of et your calculations.) a. Calculate the variance of portfollio returns, assuming the correlation between the returns is 1. (Do not round intermediate calculations. Round your answer to 4 decimal places.) b. Calculate the variance of portfolio retuns, assuming the correlation is 4. (Do not round intermediate calculations. Round your answer to 4 decimal places) Portfolio variance c. Calculate the varlance of portiolio retuns, assuming the correlation is O. (Do not round intermediate calculations. Round your answer to 4 decimal places.) Portfolo varianoe
Answer #1

a)  \sigma P =\omega I^{2}*\sigma I^{2}+\omega J^{2}\sigma J^{2}+2\omega I\sigma I\omega J\sigma J\rho

Portfolio variance = 0.452 *0.172 +0.542 *0.32 +2*0.45*0.17*0.54*0.3*1 = 0.05688225 = 5.69%

b) Portfolio Variance =  0.452 *0.172 +0.542 *0.32 +2*0.45*0.17*0.54*0.3*4 =0.13124025 =13.12%

c)    Portfolio Variance =  0.452 *0.172 +0.542 *0.32 +2*0.45*0.17*0.54*0.3*0 = 0.03209625 =3.20%